### v2

# Conceptualisation, Methodology, Software, Validation, Writing - original draft

Abstract
To correct the outdated and incomplete topology of low-voltage distribution networks (LVDNs) solely based on voltage magnitudes, a data-driven approach is developed based on machine learning algorithms and correlation analysis. Meanwhile, to address the similarity among smart meter (SM) data induced by distributed photovoltaic (PV) systems, a time-based SM data selection strategy is combined with the proposed correlation analysis. Unlike offline approaches, the proposed approach uses up-to-date voltage magnitudes to help distribution system operators determine switch states via supervised learning and refine user–feeder connections and customer phase labels through a modified hierarchical clustering algorithm. The feasibility and robustness of the proposed approach are validated using modified real-world LVDNs and multiple incomplete SM datasets collected from customers in the Netherlands. The results demonstrate that the time-based SM data selection strategy effectively mitigates its impact on phase identification, and the corrected topology not only improves network observability but also supports network operators in load balancing and PV consumption.
keywords:
[…]
Introduction
Balanced distribution network contributes to reducing voltage violations, line congestion, power loss, etc. Most existing phase balancing and reconfiguration problems are formulated as mixed-integer optimization problems that depend on network topologies [10098964]. However, these topologies are often unavailable or inaccurate for distribution system operators (DSOs) due to outdated or missing data, such as limited phase labels [vanin2024phase]. The increasing uncertainty of distributed energy resources (DERs), including household photovoltaic (PV) systems, heating pumps etc., impacts the topology reconfiguration frequency and challenges the identification of low-voltage distribution network (LVDN) topology [10026490, 10347462, 10475702]. Moreover, existing topology identification methods rely on complete smart meter (SM) datasets, the availability of which is often limited due to privacy concerns [9696306, costa2022identification]. Synthetic networks and benchmark models presented in [birchfield2016grid, 2020Non] are insufficient to represent the diversity of European LVDNs for practical use by DSOs (e.g., state estimation). Thus, practical topology correction methods for LVDNs are needed for active management and DERs penetration.
The topology identification problem in distribution networks (DNs) could be briefly classified into three categories based on the main issues they are aiming to solve: (1) switch state identification, (2) connection lines and their length identification, and (3) phase identification. The switch state identification was modeled as a classification problem in [9758816], solved by a trained deep neural network and providing quick estimations under the selected features. To recover accurate connection lines between transformers and customers, a probabilistic graphical model-based method was developed in [8456535] by restricting the propagation of SM data error. A power band-based strategy is proposed for dataset selection, aiming to enhance the accuracy of phase identification and transformer-to-meter matching [10533846]. In [liu2023hybrid], a hybrid data-driven approach integrating a partial correlation analysis strategy and a linear regression model is proposed to estimate topology from limited SM data. A similar work in [wang2023data] established a data-driven model based on the impedance matrix to correct topology and identify the impedance of connection lines from SM data.
Based on an alternating direction method of multipliers, a robust topology identification method is proposed to estimate the line connection and their parameters using $\mu$-PMUs and SM datasets [9729419]. Based on incomplete and asynchronous SM data, an optimization-based data-driven approach is constructed to recover the approximate topologies in LVDNs with open datasets [liu2025topology]. A two-step approach is introduced to identify the topology of LVDNs with multiple meter types, which provide different resolution datasets, mitigating the dependence on the SMs [10400804]. For networks with low observability, a distance matrix is introduced in [10612997] to correct the topology and is also effective in detecting latent nodes positioned between observed nodes along network paths (i.e., intermediate latent nodes rather than endpoints). A physics-informed neural network (PINN) is constructed to estimate the topology by representing the nodes without meters, whose parameters are adjusted by a reinforcement learning-based approach [10138375]. Although the above methods generate accurate topologies of DNs, the requirements of complete SM data and prior topology information (e.g., the topology candidates) make them infeasible in practice due to data privacy-related concerns. Moreover, optimizing the line resistance and reactance as independent variables may lead to unrealistic estimation, as the properties are inherently interdependent and influenced by physical environments, such as temperature, humidity.
Related to phase identification, current approaches rely on voltage magnitude data and demand profiles [7098424, garcia2025data]. Since the correlation of time-series voltage magnitude data from the same phase is stronger than that from different phases, correlation analysis is normally used. Given voltage magnitude datasets, spectral clustering and $k$-means clustering approaches are adopted to identify customer phase label [8955915]. The mixed integer linear programming model in [9409966] is constructed to identify the phase labels of single-phase customers by minimizing the gap between transformers’ measurements and the estimated data (e.g., power and voltage magnitude) in each phase. An optimization model-based approach was introduced in [10098964] to estimate customer phase labels in LVDNs with limited available recordings. Saliency analysis (SA) was integrated to extract the underlying features aiming to enhance the accuracy of phase label identification [7604110] SA and statistical tests are combined in [jimenez2020phase] to guarantee the identification accuracy of the data pairs with weak correlations. Voltage magnitude fluctuations among customers connected to different phases may exhibit similarities due to variations induced by the medium-voltage (MV) network or similar load patterns. To mitigate this effect, a high-pass filter is applied during time-series data pre-processing in the phase identification approach [hosseini2020machine]. Beyond hard classification (i.e., assigning customers a binary label of 0 or 1), soft classification is also considered for phase identification. The method in [garcia2023phase] assigns three probability values to customers. The highest probability indicates the most likely phase label for the user.
Figure 1: Deployment patterns of LVNDs: (a) feeder deployment under one street, (b) distribution of MV/LV transformers and (c) underground feeders connected to the same transformer.
Acronyms
MV/LV Medium/Low voltage LVDN Low-voltage distribution network DNs Distribution networks SM Smart meter DSO Distribution system operator DERs Distribution energy resources SA Saliency analysis PV Photovoltaic EV Electical vehicles HC Hierarchical clustering algorithm RF Random forest algorithm KNN K-nearest neighbors algorithm GNB Gaussian Naive Bayes algorithm SVM Support vector machine algorithm ANN Artificial neural network PINN Physics-informed neural network
        PCC Pearson correlation coefficient         MFP Modified PCC based Fisher z-transformation         GIS Geographic Information System         CN1 Network with two complex-structured feeders         CN2 Network with one simple- and one complex-structured feeder         SN Network with two simple-structured feeders         SNL Network with two longer simple-structured feeders         SNB Network with two simple-structured feeders with more branches         CC_PV CN1 with PV at each house         SC_PV SN with PV at each house         $C$ Complex feeder from CN1         SC Simple feeder from SN         CC_3_NU Complex feeder with three-phase non-uniform loads         SC_3_NU Simple feeder with three-phase non-uniform loads
However, the key step in the above approaches is the stepwise subtraction of the identified customers or users with the highest possibility labels, which ensures the accuracy of the phase labels of the customers who are located far away from the substation. Nevertheless, this key step might induce inaccuracies, since the accuracy of the current step is subject to that of the previous one. To address this issue, based on a modified genetic algorithm, a novel data-driven phase identification approach is introduced to identify the phase labels under incomplete datasets [jimenez2021new]. Based on multiple linear regression models, [cunha2020automated] presents a hybrid topology identification method to simultaneously identify line parameters and phase connection from raw SM data. Nevertheless, the underlying assumptions of the aforementioned methodologies—such as the presumption of complete datasets devoid of SM errors, and the omission of MV network impacts—significantly constrain their practical applicability [therrien2021assessment]. Fig. 1 presents several underground feeder deployment patterns in the Netherlands [web1], which indicates the challenges of LVDN topology identification and correction in practice, e.g., multiple cables under the same street. In addition, some of the above approaches are sensitive to missing data and measurement errors, e.g., SA-based approaches. Furthermore, the user-feeder connection identification issue is rarely considered in the aforementioned papers due to the assumption that one main feeder connects to the transformer.
To overcome these limitations, we propose a data-driven approach for topology correction in LVDNs using incomplete voltage data. Table 1 highlights the key differences between the proposed approach and existing approaches. By relying solely on voltage magnitude data, this approach preserves privacy and minimizes SM updates, as it releases the need for phase angle measurements from households. The corrected topology provides the foundational network visibility needed to implement load balancing, facilitate DERs integration, and alleviate LV line congestion, enabling active management. The main contributions are summarized as follows:
  • A data-driven topology correction approach is proposed to correct the missed and outdated recordings in LVDN topologies by three steps: (1) switch state identification, (2) user-feeder correction, and (3) phase label identification. The proposed approach relies solely on voltage magnitude data, reducing privacy risks. Compared to previous methods that depend on complete SM datasets, this approach allows the DSOs to correct the topologies using limited SM data with noise from incomplete household measurements (i.e., not all house SM is accessible).
  • The operational states of the switch bars (including multiple switchgear) are encoded using Label Encoding rules, transforming the independent identification task of each switchgear into a classification problem that can be effectively addressed using existing supervised learning methods. In correlation analysis, a modified distance function is integrated into traditional Hierarchical Clustering (HC) to amplify the distinction between loads that are geographically close but belong to different phases or cables, thereby improving identification accuracy.
  • To mitigate the impact of PV power on voltage-magnitude based correlation analysis, we propose a time-based SM data selection strategy to extract voltage magnitude measurements primarily influenced by household loads (e.g., during nighttime or low-PV-output periods). The resulting datasets enhance topology correction accuracy, particularly in scenarios with noisy or incomplete datasets.
The remainder of this paper is organised as follows: Section 2 details the framework of the proposed approach, including identification of switch states, user-feeder connections, phase labels, and voltage data pre-processing. Section 3 presents the case studies and results. Finally, Section 4 summarizes the conclusions of this work.
Figure 2: Proposed topology correction approach composed of three steps: (1) switch state identification, (2) user-feeder connection identification and (3) phase labels identification.
Table 1: Input data and outputs in topology problems
Method Ref. Input Data Robustness Output
$V$ $P/I$ Limited GIS DERs SM Missed Switch user-feeder Phase
topology data impact errors data state connections labels
DNN [9758816] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
[10138375] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
Linear regression [8456535] $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
[cunha2020automated] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Optimization based [liu2025topology] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
[9409966] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
clustering-based [8955915] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
[10533846] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Heuristic Algorithms [jimenez2021new] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
Our Method $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Topology Correction Approach
The proposed topology correction approach is illustrated in Fig. 2, which is composed of three steps: (1) switch state identification, (2) user-feeder connection identification and (3) phase label identification. The input includes customer voltage magnitude and incomplete recordings of the user-feeder connection, and the output is the corrected topology with accurate customer’s user feeder connections and phase labels, including single-phase and three-phase customers.
Switch State Identification
In LVDNs, multiple switch bars exist, each containing a different number of switches. Each switch connects two cables (e.g., A and B), which are linked to different transformers (e.g., AT and BT). Under normal conditions, the switch remains open. When transformer AT becomes overloaded, the switch is closed, shifting the load on cable A to transformer BT. Therefore, the switch state plays a crucial role in maintaining network efficiency and enabling effective load shifting. An example of switch deployment is illustrated in Fig. 3. This non-uniform deployment poses challenges for accurately identifying the state of individual switches using SM data. To address this, a random forest (RF) algorithm is employed, leveraging voltage magnitude datasets to determine the state of switches within each switch bar. First, inspired by Label Encoding rules, the switch states in one bar are encoded into single-digital label formats, i.e., each combination of the switch states in the switch bar is uniquely mapped to a single digital label. The number of switches is assumed to be accessible for the DSO. For example, when there are 3 switches in the switch bar, the switch states $SS$ can be encoded as follows:
$f(V)=\begin{cases}0,&\hskip 22.76228pt\text{if}\quad SS=[1,1,0]\\ 1,&\hskip 22.76228pt\text{if}\quad SS=[1,0,1]\\ 2,&\hskip 22.76228pt\text{if}\quad SS=[0,1,1]\end{cases}$ (1)
After the switch states are encoded, the switch state identification problem in each bar is transformed into a standard multi-class classification task. The input for this classification task consists of time-series voltage magnitude datasets. However, the presence of missing data from SM or unmetered households results in varying numbers of available SMs on each feeder. Feeders with a higher number of available SM provide more data for the supervised learning process, potentially introducing bias in classification and leading to accuracy loss. Besides, when switch states remain unchanged for extended periods (e.g., a state $[1,0,1]$ resulting in an excessive number of samples labelled as $1$), the dataset becomes unbalanced. This imbalance further affects classification accuracy, posing a common challenge in machine learning.
Figure 3: Illustrative example of switch deployment and states in LVDNs: 1 represents the activated switch and vice versa [abb2025].
To address the above-mentioned challenges, an SM selection process is employed before switch state identification. This selection process consists of two key steps:
  • Uniform SM number selection across feeders. $M$ represents the minimum number of available SMs across all feeders. To ensure consistency, the number of SMs used for data extraction from each feeder is fixed at $M$. Specifically, the $M$ SMs closest to the switch bar are selected as representative features of that feeder.
    $M=\min\{m_{1},m_{2},\dots,m_{K}\},$ (2)
    where $m_{i}$ is the number of available SMs on feeder $i$, and $K$ is the total number of feeders that are connected to the same switch bar.
  • Balanced dataset construction. Given the encoded switch state labels (i.e., 0, 1, 2, …), an equal number of samples is selected for each category. For each encoded class $C_{i}$, the same number of samples $S$ is randomly selected.
    $S=\min\{|C_{1}|,|C_{2}|,\dots,|C_{J}|\},$ (3)
    where, $|C_{j}|$ represents the number of available samples for class $j$, and $J$ is the total number of classes.
Through this selection process, the impact of the varying number of available SMs and class imbalance is minimized, leading to a more balanced and effective training dataset for the multi-class classification task. The selected voltage measurements can be represented in a matrix form $V_{train}$, forming the input for the multi-class classification task. Each row in matrix $V_{train}$ represents training one sample with an encoded $SS$ label. $v^{s,C_{j}}_{m_{k},M}$ represents the voltage magnitude of the house $M$ connected to the feeder with $m_{k}$ houses when the switch state belongs to category $C_{j}$.
$V_{train}=\begin{bmatrix}v^{1,C_{1}}_{m_{1},1}&\cdots&v^{1,C_{1}}_{m_{1},M}&\cdots&v^{1,C_{1}}_{m_{k},1}&\cdots&v^{1,C_{1}}_{m_{k},M}\\ \vdots&&\vdots&&\vdots&&\vdots\\ v^{S,C_{1}}_{m_{1},1}&\cdots&v^{S,C_{1}}_{m_{1},M}&\cdots&v^{S,C_{1}}_{m_{k},1}&\cdots&v^{S,C_{1}}_{m_{k},M}\\ \vdots&&\vdots&&\vdots&&\vdots\\ v^{1,C_{J}}_{m_{1},1}&\cdots&v^{1,C_{J}}_{m_{1},M}&\cdots&v^{1,C_{J}}_{m_{k},1}&\cdots&v^{1,C_{J}}_{m_{k},M}\\ \vdots&&\vdots&&\vdots&&\vdots\\ v^{S,C_{J}}_{m_{1},1}&\cdots&v^{S,C_{J}}_{m_{1},M}&\cdots&v^{S,C_{J}}_{m_{k},1}&\cdots&v^{S,C_{J}}_{m_{k},M}\\ \end{bmatrix}$ (4)
Figure 4: A practical LVDN with multiple parallel feeders along the same street demonstrates the common challenge of ambiguous use-feeder connections and phase labelling, illustrating the physical complexity that makes accurate network topology identification (i.e., step (2) and (3)) difficult.
The RF algorithm excels in multi-class classification by handling high-dimensional data, capturing complex feature interactions, and maintaining accuracy despite noise. Its ensemble learning approach—building diverse trees with random feature selection—prevents overfitting and enhances generalisation. RF also provides feature importance insights, making it ideal for switch state identification. Given a dataset $D=\{(X_{1},y_{1}),\ldots,(X_{n},y_{n})\}$ where $X_{i}$ represents the feature vector (i.e., each row in $V_{train}$) and $y_{i}$ is the encoded label by (1), the general process of RF training read as:
  1. 1.
    Draw $B$ bootstrap samples from the training data $V_{train}$ and the corresponding labels.
  2. 2.
    Grow an unpruned decision tree for each sample by randomly selecting a subset of features at each split.
  3. 3.
    For classification, each tree votes on the most common class label, and the RF outputs the majority vote:
    $\hat{y}=\text{argmax}_{j}\sum_{b=1}^{B}\mathbb{I}(h_{b}(x)=j),$ (5)
where $h_{b}(x)$ is the prediction of the $b$-th tree and $\mathbb{I}$ is the indicator function. Given a voltage sample $V_{i}$, each trained decision tree independently predicts $h_{b}(x)$ and the final labels of the switch are obtained by the majority voting mechanism.
User-feeder Connections Identification
The connection relations between customers and the feeders might be outdated and not always available for DSOs, which is ignored by multiple similar works [liu2025topology, 10138375], especially for houses and multiple-layer buildings that are close to the streets with multiple deployed underground feeders. A practical case is shown in Fig. 4. Those unclear user-feeder connections also impact the phase label identifications. A correlation approach is introduced to reveal the user-feeder connections, which only rely on time-series voltage magnitude datasets. Two practical cases are addressed: 1) user-feeder connection identification without reference and 2) user-feeder connection identification with limited, accurate recordings.
Connection Identification Without Recordings
The proposed user-feeder connection identification algorithm combines the Hierarchical clustering (HC) algorithm and the correlation analysis based on the modified Pearson correlation coefficient (PCC). Algorithm 1 depicts the process of the proposed approach. The voltage magnitude dataset is stored in a matrix $V$, which is used as the input for the algorithm. Vector $V_{n}=[v_{n,1},v_{n,2},...,v_{n,\cal{T}}]^{T}$ represents the time series voltage magnitude at household $n$. $v_{n,t}$ represents the voltage magnitude at household $n$ at time $t$. $T$ is the measurement period of the SM data.
Input: $V$, $K_{s}$, NN
for $n\leq N$ do
 for $m\leq N$ do
    $\text{PCC}(V_{n},V_{m})$ via Eq. (6)
    $\text{FPCC}(\rho_{n,m})$ via Eq. (7)
   end for
end for
for $i\leftarrow 1$ to NN do
 for $j\leftarrow i+1$ to NN do
    $D_{ij}\leftarrow\text{MFP}(V_{n},V_{m})$ via Eq. (8)
   end for
end for
Obtain $\mathbf{D}\leftarrow\{D_{ij}\}$
Clusters $\mathcal{C}$: $\leftarrow$ Traditional HC and $\mathbf{D}$
labels $\leftarrow$ Assigned labels to households based on $\mathcal{C}$
Output: labels, Cluster: $C_{1}$, …, $C_{K_{s}}$
Algorithm 1 MFP-based User-Feeder Identification without Available Labels
$PCC(V_{n},V_{m})=\frac{\sum_{t=1}^{T}(v_{n,t}-\bar{V_{n}})(v_{m,t}-\bar{V_{m}})}{\sqrt{\sum_{t=1}^{T}(v_{n,t}-\bar{V_{n}})^{2}}\sqrt{\sum_{t=1}^{T}(v_{m,t}-\bar{V_{m}})^{2}}}\quad$ (6)
where $\bar{V_{n}}$ and $\bar{V_{m}}$ are the mean values of samples $V_{n}$ and $V_{m}$, respectively.
Usually, the PCC of voltage magnitudes in the same feeders shows higher values compared to those in different feeders [7098424]. To amplify the difference between voltage correlation from different feeders, a modified Fisher $z$-transformation function is employed:
$FPCC(V_{n},V_{m})=\ln[\frac{1+PCC(V_{n},V_{m})}{1-PCC(V_{n},V_{m})+\alpha}],$ (7)
where $\alpha$ added to the denominator is used to avoid an infinite value of $FPCC(\cdot)$ and to control its distribution region.
To amplify the difference between voltage from different feeders, a modified $FPCC(\cdot)$ value (MFP) is employed to replace the Euclidean distance in the traditional HC algorithm.
$MFP(V_{n},V_{m})=1-min\{\frac{1}{4}\ln(1+e^{a+4\cdot FPCC(V_{n},V_{m})}),1\}$ (8)
The second term in (8) is a modified likelihood function $F(\cdot)$ [garcia2023phase], which is used to calculate the $MFP(\cdot)$ between each sample $V_{n}$ and the rest of the samples. Since the range of the function $F(\cdot)$ is $(0,1]$, the range of the distance metric $MFP(\cdot)$ is correspondingly $[0,1)$. A value of $MFP(V_{n},V_{m})$ closer to 0 indicates a higher likelihood that the two samples $V_{n}$ and $V_{m}$ are collected from the same substation.
Given the switch state obtained from Section 2.1, the feeder-transformer connection is available, and the number of feeders connected to the same transformer is assumed to be accessible for the DSO and taken as the number of clusters (i.e., $K_{s}$). Lines 1-13 in Algorithm 1 are to obtain the MFP of the input voltage magnitude dataset. Lines 16-23 are to find the link matrix of matrix $D$. From line 23 onwards, input data are clustered into a $K_{s}$ cluster using the obtained linkage matrix. If two users share the same label, they are connected to the same underground feeder, and vice versa.
Input: $V$, $N_{k}$, $N_{un}$, $L_{un}$, $k$
for $n\geq N_{un}$ do
 for $m\leq N_{k}$ do
    $_{0}(n,m) \leftarrow DD$($V_{n}$$V_{m}$) via Eq. (8)
   end for
 $\hat{y}=\arg\max_{c}\sum_{n\in\mathcal{N}_{k}}\mathbb{I}(y_{n}=c)$
 $L_{un}[n]$ $\leftarrow$ $\hat{y}$
end for
Output: $L_{un}$ for households with unknown labels
Algorithm 2 MFP-based User-Feeder Identification with Available Labels
Connection Identification With Recordings
When limited user-feeder connections are known, the user-feeder identification problem becomes more manageable compared to the fully unknown scenario. There are two sources of accurate recordings:
  • Historical recordings from DSO’s datasets.
  • Identified connections with high confidence from Section  2.2.1.
Those available recordings transform the identification problem from an unsupervised learning problem to a supervised learning problem. Various supervised learning algorithms could be used to identify unknown connections, e.g., ANN, KNN, etc. The traditional KNN algorithm and the simplified MFP-based approach were employed and compared. The KNN algorithm is a widely used supervised learning method for classification tasks. It classifies an unknown sample by considering the labels of its nearest neighbours in the feature space, which is similar to identifying the user-feeder connection based on their location and the neighbours with known labels. Given a set of labelled training samples, the classification of a new sample is determined based on the majority class among its $k$ nearest neighbours. The distance between samples is commonly measured using the Euclidean distance
$d(V_{n},V_{m})=\sqrt{\sum_{t=1}^{T}(v_{n,t}-v_{m,t})^{2}}.$ (9)
Once the distances between the new sample and all labelled samples are computed, the $k$ nearest neighbours are selected. The final classification is determined using majority voting.
$\hat{y}=\arg\max_{c}\sum_{n\in\mathcal{N}_{k}}\mathbb{I}(y_{n}=c)$ (10)
$\mathcal{N}_{k}$ denotes the set of $k$ nearest neighbors, $y_{i}$ represents the label of neighbour $i$. In user-feeder connection identification, the KNN algorithm assigns users without clear connections to the most probable feeder based on similarity to datasets from users who have available user-feeder connection recordings. Motivated by the traditional KNN approach, the proposed MFP-based approach is simplified to reveal the unknown user-feeder connection in DNs, which is depicted in Algorithm 2. $\mathcal{N}_{un}$ is the number of user with unclear connections, and $\mathcal{L}_{un}$ is their identified labels set.
Phase Label Identification
Figure 5: Illustrative example showing the sequential process and intermediate outputs of the proposed topology correction framework, demonstrating (1) switch states, (2) user-feeder connections and (3) phase labels.
Figure 6: Distribution of voltage magnitudes in a LVDN:(a) winter datasets and (b) summer datasets.
After the second step in the proposed framework (i.e., user-feeder connection identification in Section 2.2), the third step is to identify the phase labels of the customer connection for each feeder, which is a typical 3-cluster classification problem. The phase identification approach is introduced by modifying Algorithm 1, i.e., removing Fisher $z$-transformation and using normalized voltage data $V^{*}$ as inputs.
$V=\begin{bmatrix}v_{1,1}&\cdots&v_{1,t}&\cdots&v_{1,T}\\ v_{2,1}&\cdots&v_{2,t}&\cdots&v_{2,T}\\ \vdots&&\vdots&&\vdots\\ v_{N,1}&\cdots&v_{N,t}&\cdots&v_{N,T}\end{bmatrix}\hskip 28.45274pt$ (11)
$v_{n,t}^{*}=\frac{v_{n,t}-\mu_{t}}{\sigma_{t}},\hskip 71.13188ptn\in\mathcal{N},t\in\mathcal{T}$ (12)
$v_{n,t}$ is the value of the $t$-th feature for the $n$-th user. $\mu_{t}$ is the mean of the $t$-th feature, and $\sigma_{t}$ is the standard deviation of the $t$-th feature. Given the normalized voltage dataset $V^{*}$ and set $K_{s}$ as 3, modified Algorithm 1 is adopted to identify the phase labels of each house. The output of each step is depicted in Fig. 5, and the obtained topology presents the clear connections between users and transformers feeders, and phases.
Time-based Data Selection
Compared to traditional LVDNs, the voltage magnitude measurements in modern LVDNs may exhibit higher values due to the integration of PV systems. Within the same distribution network area, environmental factors such as light intensity, temperature, and humidity remain relatively uniform, leading to similar PV generation patterns among users. The voltage magnitudes from the same LVDN are depicted in Fig. 6. This effect is particularly pronounced during summer when high-power PV systems are connected (i.e., the voltage magnitude distribution in Fig. 6(b)). Consequently, users with identical PV capacities but connected to different phases may experience analogous voltage magnitude fluctuations. This phenomenon introduces challenges to the identification steps based on correlation analysis, as the voltage magnitude variations induced by PV generation may obscure the underlying phase-dependent characteristics.
Although various signal processing techniques, such as SA and high-pass filters, can be employed to mitigate high-frequency or low-frequency noise induced by household PV, the determination of appropriate parameters—such as cutoff frequency—must be tailored to specific datasets. This requirement presents practical challenges in implementing these methods effectively. Inspired by the power-band based strategies in [10533846], a time-based data selection strategy is introduced to enhance identification accuracy by screening summer data, especially for phase identification. The proposed time-based data selection strategy is easily deployed and fast. Two strategies for selecting relevant data are introduced:
  • Fixed time threshold. Data is selected from a predetermined time window, such as from 22:00 to 5:00, when PV generation is minimal.
    $\mathcal{T}_{d}=\{t\mid t_{1}\leq t\leq t_{2}\},$ (13)
    where $t_{1}$ and $t_{2}$ define the fixed time window within which the data are extracted.
  • Dynamic time threshold. Data is selected based on historical PV output values, ensuring that the chosen data corresponds to periods when PV generation is lower than the user load.
    $\mathcal{T}_{d}=\{t\mid P_{\text{PV}}(t)<P_{\text{load}}(t)\},$ (14)
    where $\mathcal{T}_{d}$ represents the selected time instances, $P_{\text{PV}}(t)$ is the PV output at time $t$, and $P_{\text{load}}(t)$ is the user load at time $t$. Only the data points satisfying this condition are retained for phase identification.
By applying these time based selection strategies, the impact of PV-induced voltage fluctuations can be mitigated, thereby improving the robustness of correlation-based phase identification methods.
Case Study
In this section, the feasibility and robustness of the proposed topology correction framework are validated using modified real-world LVDNs and SM data collected from customers in the Netherlands. The code used in this paper is available online  [^1] https://github.com/distributionnetworksTUDelft/Data-driven-approach-for-topology-correction-in-LVDNs-. The resolution of the dataset is 15 minutes, resulting in 96 data points per day. The base three-phase voltage level is 0.4 kV, and the loads of three-phase customers are non-uniformly distributed across Phases A, B, and C. The time thresholds $t_{1}$ and $t_{2}$ are set to 20 and 88. The voltage magnitude profiles are generated using a power flow model, being solved by using the Power Grid Model package [PowerGridModel].
Switch State Analysis
Figure 7: Switch state identification accuracy under incomplete dataset by: a) RF and b) SVM
Table 2: Accuracy (%) comparison under multiple scenarios
Number Number Methods
of Switch of available SM RF SVM GNB
3 2 69.16 63.31 60.39
10 87.98 88.63 85.71
20 88.31 88.96 84.74
All 95.45 96.75 89.61
5 2 56.56 53.59 56.72
10 91.56 99.38 95.16
20 92.50 99.21 94.06
All 98.75 98.59 95.94
The switch bar is connected to three and five transformers, respectively, with all transformers assumed to be connected to the same MV network, which introduces similar low-frequency variations in customer voltage profiles. Each feeder connects to 49 customers. Using the voltage datasets selected through the time-based strategy described in Section 2.4, the performance of three models is evaluated under varying numbers of available SMs at each feeder, including RF, Support Vector Machine (SVM), and Gaussian Naive Bayes (GNB).
Table 2 demonstrates that the accuracy of all methods improves as the number of available SM increases. SVM achieves high accuracy (i.e., above 98% for 5 switches scenarios and 88% for 3 switches scenarios) once the number of houses reaches 10, while RF reaches the same level only when all SM data are available. This suggests that SVM is more efficient in learning from limited data, likely due to its ability to find optimal decision boundaries even with fewer samples. For the two SM data available scenarios, the accuracy of RF is about 5% higher than that of SVM in two scenarios (i.e., 3 and 5 switches in the switch bar). Moreover, GNB underperforms compared to RF and SVM, with its accuracy lagging behind when the number of houses is low (lower than 90% and 95% under all the SMs data in two cases, respectively).
Given datasets with random missed data points, the accuracy of RF and SVM were assessed, and the average accuracy under five simulations is shown in Fig. 7. In general, the accuracy of the two approaches improves significantly as the number of available SMs increases and slightly decreases as the missing percentage increases, demonstrating the robustness of the proposed method with sufficient data availability. When only 2 SMs data are available, the accuracy remains consistently low (around 55–60%), indicating the method’s limited capability under sparse data conditions. Besides, the sensitivity of SVM accuracy to the available data is less than that of RF, since the accuracy of SVM remains close when the number of available SM is larger than 10. These findings highlight the robustness of SVM and RF for switch state identification tasks in LVDNs.
Figure 8: Topology of CN2 in (a) and SNB in (b).
Table 3: Accuracy (%) comparison under multiple scenarios
Network Phase Per-node Two Three
Acronyms connection feeders feeders
CN1 1 multiple 100 100
single 100 96.42
3 multiple 67.44 55.04
single 50.00 46.43
CN2 1 multiple 100 98.26
single 100 100
3 multiple 42.86 30.43
single 62.50 35.56
SN 1 multiple 94.56 67.56
single 75.56 38.46
3 multiple 55.56 40.54
single 48.89 24.15
SNL 1 multiple 96.30 97.33
single 82.22 84.84
3 multiple 59.26 37.33
single 46.67 30.30
SNB 1 multiple 95.45 92.50
single 80.00 88.57
3 multiple 55.56 43.18
single 48.49 45.71
User-feeder Identification Analysis
The influence of network complexity, phase configuration, and per-node connection on classification accuracy was first analysed. The per-node connection is the number of households connected to a point, e.g., a multi-unit building. Given datasets with SM errors but without available recordings of user-feeder connections, Table 3 presents the accuracy comparison under five modified configurations of LVDNs: a network with two complex-structured feeders (CN1), a network with a simple-structured and one complex-structured feeder (CN2), a network with two simple-structured feeders (SN), network with longer simple-structured feeders (SNL), and a network with two simple-structured feeders with more branches (SNB). Another complex-structured feeder or simple-structured feeder is added to the above five LVDNs to construct LVDNs with three feeders (as shown in the fifth column). The main difference between a complex network and a simple network is the length of feeders and the number of branches. The topology of CN2 and SNB are shown in Fig. 8.
Figure 9: PCC in first row and FPCC in second row among voltage datasets: a) with single-phase customers and b) with three-phase customers
Table 4: Average accuracy under three-phase datasets and multiple connections at nodes with SM errors
Network Unknown Modified KNN Traditional KNN
Acronyms rate $k=1$ $k=3$ $k=5$ $k=8$ $k=1$ $k=3$ $k=5$ $k=8$
CN1 10% 98.16% 98.10% 99.02% 99.59% 99.90% 99.79% 99.84% 99.95%
20% 96.16% 96.49% 97.44% 98.94% 99.84% 99.67% 99.68% 99.51%
30% 93.73% 95.19% 96.73% 98.65% 99.52% 99.48% 99.08% 98.13%
40% 91.71% 92.79% 96.13% 97.84% 99.32% 98.75% 97.92% 94.70%
50% 88.84% 90.63% 95.59% 96.92% 99.03% 98.33% 94.76% 88.51%
SN 10% 97.63% 98.06% 98.41% 98.57% 99.65% 99.65% 99.20% 98.31%
20% 95.07% 95.80% 97.04% 97.11% 99.24% 99.30% 98.24% 96.98%
30% 92.19% 93.80% 95.20% 94.72% 98.98% 98.20% 97.00% 92.22%
40% 90.11% 91.37% 92.50% 92.44% 98.56% 97.48% 93.54% 88.94%
50% 87.11% 88.70% 89.43% 87.43% 97.72% 92.98% 88.39% 83.50%
SNB 10% 97.43% 98.09% 98.33% 98.43% 99.61% 99.81% 99.48% 98.65%
20% 94.37% 95.87% 96.56% 96.83% 99.39% 99.31% 98.63% 96.52%
30% 91.80% 93.19% 94.78% 94.83% 98.81% 98.43% 96.69% 92.54%
40% 90.35% 91.70% 93.02% 91.87% 98.43% 96.78% 93.57% 89.04%
50% 87.50% 88.46% 89.31% 87.22% 96.76% 94.15% 88.57% 82.11%
Identification Without Recordings
The PCC and MFP among customers in CN1 with multiple connections per node are depicted in Fig. 9. As illustrated in Fig. 9(a), the expression in (7) reduces the voltage correlation between users connected to different feeders. This step helps distinguish between different connections more clearly. However, the complexity of the voltage data correlation shown in Fig. 9(b)—reflected in the similarity of the correlation distribution — makes it extremely challenging to simultaneously identify both the phase label and the user-feeder connection using only voltage data. A balanced LVDN represents a scenario where all users are connected to the same phase (i.e., a single-phase LVDN), whereas an unbalanced LVDN corresponds to a three-phase configuration.
Table 3 summarizes the accuracy of user-feeder connection identification across five LVDN scenarios. In general, the proposed approach demonstrates higher robustness to variations in the number of feeders connected to the same transformer and per-node connections but exhibits lower robustness in three-phase configurations. Specifically, accuracy drops from 100% to 96.43% in CN1 with three feeders and multiple connections per node, and from 94.56% to 55.56% in SN with two feeders and multiple connections per node, highlighting the impact of complex multi-phase settings. In simpler LVDNs, per-node connections significantly affect accuracy. For instance, in SNL with two feeders in a single-phase setting, accuracy declines from 96.3% to 82.22%, while in SNL with two feeders in a three-phase setting, it drops from 59.26% to 46.67%. In contrast, for complex LVDNs, accuracy remains above 95% regardless of single or multiple per-node connections, indicating greater resilience in more intricate network structures. Moreover, accuracy in complex LVDNs is slightly higher than in simpler LVDNs, primarily due to feeder length differences. Simple LVDNs, which are modified from urban areas with high population density, have houses located close to each other (i.e., smaller distances between users with different connections). This proximity reduces the difference in the voltage variations between neighbouring houses, leading to greater identification accuracy. However, the length of the feeder and the number of branches have a minor effect on the accuracy, which is verified in the three simple LVDNs (i.e., SN, SNL, and SNB).
Since unbalanced LVDNs are more representative of real-world conditions compared to balanced LVDNs, only half of the identified user-feeder connections with high confidence are selected for further processing in the subsequent section. These experimental results highlight the inherent challenge of simultaneously identifying user-feeder connections and accurately determining the phase labels of individual users. The complexity arises from the intertwined nature of voltage variations across different feeders and phases, further complicated by factors such as DERs and seasonal load fluctuations.
Identification With Recordings
Given datasets with varying levels of unknown user-feeder connections, Table 4 compares the average accuracy of 20 simulations at each scenario (i.e., each unknown level). The results provide insights into the robustness and effectiveness of both methods across different network configurations. In general, as the percentage of unknown user-feeder connections increases from 10% to 50%, a gradual decline in accuracy is observed across all methods and network configurations. However, both the proposed MFP-based method and KNN (with varying $k$ values) maintain relatively high performance, with accuracy exceeding 91% even at 30% unknown connections.
Table 4 indicates that the traditional KNN algorithm consistently outperforms the proposed modified KNN approach when the parameter $k$ is less than 8. However, the modified KNN demonstrates improved robustness as the proportion of missing data increases, particularly at higher unknown rates and a larger $k$. Specifically, the modified KNN method outperforms traditional KNN at higher levels of unknown information (e.g., 50%), achieving accuracy levels exceeding 87%. This suggests the proposed method may better preserve identification accuracy under limited accurate recordings from DSOs. In the simple network with branches (SNB), the performance gap between the two methods becomes less pronounced. Traditional KNN shows a slight advantage at intermediate unknown rates (e.g., 15%–20%) and smaller $k$ values, whereas the modified KNN achieves comparable or superior performance at higher levels of missing data and larger $k$ values. Fig. 10 depicts the accuracy of the approaches under datasets with various missing data percentages. The accuracy declines slightly by about 5% as the percentage of missing data increases, even at 70% missing data. This demonstrates the strong robustness of the proposed method against missing data.
Figure 10: Accuracy under incomplete dataset: a) modified KNN with 20% unknown labels, b) traditional KNN with 20% unknown label, c) modified KNN with 50% unknown labels, d) traditional KNN with 50% unknown labels.
Phase Identification Analysis
The influence of feeder complexity, environmental conditions, and load distribution on identification accuracy is analyzed, with results summarized in Table 5 and 6. CC_PV: Represents a complex-structured feeder from CN1, with photovoltaics (PV) installed at each house. SC_PV: Corresponds to a simple-structured feeder from SN, also with PV at each house. CC and SC: Represent feeders from CN1 and SN with PV, but with datasets preprocessed using the method described in Section 2.4. These four feeders serve single-phase users only.CC_3_NU and SC_3_NU: Denote complex and simple feeders equipped with three-phase meters, where the load is non-uniformly distributed among the three phases. “+20” indicates an additional 20 customers connected to the feeder. “Long feeder” refers to an extended feeder length compared to the original configuration. Accuracy and clustering purity are used to evaluate the performance of phase identification, as formulated in the expression (15).
$P_{pu}=\frac{1}{N}\sum_{\mathcal{L}_{k}\in{\mathcal{L}}}\underset{i}{\operatorname{max}}|\mathcal{L}_{k}\cap\mathcal{L}_{i}|\$(15)
where $\mathcal{L}$ denotes the set of obtained labels for all input data, and $\mathcal{L}_{k}$ represents the label assigned to the data in cluster $\mathcal{C}_{k}$. $\mathcal{L}_{i}$ is the label set of data points that truly belong to class $\mathcal{C}_{i}$.
Table 5: Average purity (%) under datasets form feeders with different components
Network Setting Winter Dataset Summer Dataset
single multiple single multiple
CC_PV - 100 100 78.57 65.12
+20 100 100 77.08 77.78
SC_PV - 96.00 97.06 84.00 88.24
longer feeder 100 94.12 84.00 88.24
$C$ - 100 100 100 97.67
+20 100 100 100 100
SC - 88.88 88.24 68 82.35
longer feeder 100 100 68.00 82.35
CC_3_NU PV 100 100 79.41 83.67
filtered 100 100 100 100
SC_3_NU PV 97.14 97.92 71.43 100
filtered 97.92 94.29 100 95.83
Table 6: Average accuracy (%) under datasets form feeders with different components.
feeder setting Winter Dataset Summer Dataset
single multiple single multiple
CC_PV - 100 100 88.89 72.08
+20 100 100 76.11 87.88
SC_PV - 96.97 97.78 87.88 90.30
longer feeder 100 94.75 87.88 90.30
$C$ - 100 100 100 97.78
+20 100 100 100 100
SC - 87.45 87.88 71.86 79.22
longer feeder 100 100 71.86 79.22
CC_3_NU PV 100 100 86.11 85.19
filtered 100 100 100 100
SC_3_NU PV 97.78 97.62 76.97 100
filtered 97.62 95.56 100 95.24
Overall, Table 5 indicates consistently higher accuracy in winter datasets compared to summer datasets. This discrepancy is primarily due to voltage variations caused by household photovoltaic (PV) systems, which negatively impact phase identification performance. In the first four cases, where only single-phase or balanced three-phase customers are considered, the influence of PV-induced voltage fluctuations is evident. For instance, in the summer dataset, the CC_PV feeder (a complex feeder from CN1 without PV filtering) achieves 88.89% accuracy with a single connection per node but drops to 72.08% with multiple connections per node. However, when PV-induced variations are filtered out—as in the CC and SC cases—accuracy consistently exceeds 95%, highlighting the critical role of PV filtering in improving phase identification.
These findings demonstrate that phase identification accuracy is highly dependent on seasonal variations and dataset complexity, particularly in the presence of PV-induced voltage fluctuations. Without applying the proposed time-based selection process to mitigate PV impacts, accuracy remains below 90% in most cases. In contrast, once PV-induced variations are filtered, accuracy consistently exceeds 95%, regardless of network complexity or three-phase load distribution. Additionally, factors such as the number of connections per node, feeder length, and the number of connected houses have a relatively minor impact on classification accuracy compared to PV output power fluctuations. This underscores the importance of addressing PV-induced voltage variations to achieve robust and reliable phase identification.
Conclusion
In this paper, we proposed a practical topology correction framework for LVDNs, incorporating a time-based smart meter selection strategy to mitigate the impact of PV systems and a balanced dataset construction to improve training effectiveness. The experimental results demonstrate that the proposed framework is highly robust, maintaining high accuracy even with incomplete datasets, regardless of random missing data, data dimensionality, or topological complexity. Specifically, the correlation-based user and phase identification shows strong resilience to data loss. However, the switch state identification component, based on supervised learning, is more sensitive to missing and asynchronous data. Moreover, the framework’s performance is influenced by seasonal variations and voltage fluctuations caused by PVs, highlighting the importance of time-based data selection in ensuring reliable topology identification. Future extensions of this work will account for the influence of various flexible assets to enhance the scalability and precision of the framework.
[…]